A large bakery buys flour in 25-pound bags. The bakery uses 1500 bags per
year. Preparing an order and receiving a shipment of flour involved a cost of $10 per order.
Annual carrying cost is $75 per bag.
a) Determine the optimal order quantity.
b) What is the average number of bags on hand?
c) How many orders per year will there be?
d) Calculate the total annual cost of ordering and holding flour for the EOQ value.
e) If the holding cost were to increase by $9 per year, how much would that affect the
minimum total annual cost?
Answer #1

1. DEMAND = 1500

ORDERING COST = 10

HOLDING COST = 75

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST) = SQRT(2 * 1500 * 10 / 75) = 20

2. AVERAGE INVENTORY = EOQ / 2 = 20 / 2 = 10

3. ORDER FREQUENCY = DEMAND / Q = 1500 / 20 = 75

4. ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST = (20 / 2) * 75 = 750

ANNUAL ORDERING COST = (DEMAND / EOQ) * ORDERING COST = (1500 / 20) * 10 = 750

TOTAL COST OF MANAGING = AHC + AOC = 750 + 750 = 1500

5. WITH INCREASED HOLDING COST TO 84:

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST) = SQRT(2 * 1500 * 10 / 84) = 19

ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST = (19 / 2) * 84 = 798

ANNUAL ORDERING COST = (DEMAND / EOQ) * ORDERING COST = (1500 / 19) * 10 = 789

TOTAL COST OF MANAGING = AHC + AOC = 798 + 789 = 1587

DIFFERENCE = 1587 - 1500 = 87 OR INCREASE BY $87